Calibration/Production Update

Speaker : Gene Van Buren ( BNL )


Talk time : 11:15, Duration : 00:15
Productions:
  • Weekly status is here: You do not have access to view this node
  • Isobar Phase II production completed: ~55k picoDsts on /star/data97/reco covering all 1430 runs that will be used for blind analyses
  • Run 17 pp510 st_W stream re-production with proper GridLeak distortion corrections completed:
    get_file_list.pl -keys path,filename -cond production=P20ic,trgsetupname=pp500_production_2017,filetype=daq_reco_muDst,filename~st_W,storage=nfs -limit 0
  • MuDst=>Pico conversion of Run 16 AuAu200 over 1/3rd completed
  • We will need to continue clearing out NFS space to make room for ongoing productions (this is not a case of 100s of TBs sitting around free)
    • Datasets aren't automatically moved to distributed disk and a request to do so is needed!

Calibrations:
  • Run 19 AuAu19.6:
    • TPC super-sector alignment and SpaceCharge calibrations were iterated an additional time to be sure of convergence.
    • Large calibration production then proceeded and completed for purposes of...
      • Checking BTOF & VPD calibrations
      • Checking TPC dE/dx calibrations
      • ETOF calibrations with time granularity sufficient to handle observed variations
        • Estimate was that this could take 2-3+ weeks, and can be applied in a later MuDst=>PicoDst conversion if for some reason we didn't want to wait for it
    • Physics production should start in mid-July!
  • Run 19 AuAu 7.7, 14.5, 200:
    • TPC SpaceCharge & GridLeak is done, ready for database upload today
    • Large calibration production can be done whenever we're ready to prioritize it
  • Runs 19 & 20 FXT:
    • TPC T0s?
    • Need TOF and TPC dE/dx calibrations
  • Run 20 AuAu9.2:
    • Booster Main Magnet operations will lead to exclusion of some data at the beginning of many runs (total fraction not yet well known)
    • TPC SpaceCharge & GridLeak stalled for a couple weeks [workforce has to focus on thesis]
      • Will need to check that calibration before shutdown matches the recent data from after shutdown
    • Need TOF and TPC dE/dx calibrations
  • Run 20 AuAu11.5:
    • Similar status as AuAu9.2, but will be worked on last

-Gene